calculate CEC and show the calculation (for both soil A and soil B) used to derive CEC from the measured variables. You should also calculate the acid saturation, base saturation and the exchangeable sodium percentage (ESP)


soil samples for soil (A) 157.72 K(mg/kg) 756.56 Ca(mg/kg) 198.78 Mg(mg/kg

21.36 Na (mg/kg 0.06 Al+H (cmolc/kg)


soil( B)  142.66 K (mg/kg)  1139.54 Ca(mg/kg)  157.22 Mg(mg/kg  35.89  Na (mg/kg)

0.08 Al+H (cmolc/kg)







1
Expert's answer
2021-06-25T01:28:14-0400

CEC=base cations +acid cations

for soil A

base cations= (157.72/200) +(756.56/120)+(198.78/390)+(21.36/230)=8.5879

acid cations= 0.06

therefore CEC=8.5879+0.06=8.6479 meq/100 g


for soil B

base cations= (142.66/200)+(1139.54/120)+(157.22/390)+(35.89/230)=10.7685

acid cations= 0.08

therefore CEC=10.7685+0.08=10.8485 meq/100 g



base saturation =(base cations/CEC)*100

for soil A

base saturation=(8.5879/8.6479)*100=99.31%

acid saturation =(0.06/8.5879)*100= 0.699%

for soil B

base saturation= (10.7685/10.8485)*100=99.26%

acid saturation=(0.08/10.8485)*100=0.7374%


exchangeable sodium percentage

for soil A

ESP= (0.0929/10.8485)*100=1.074%

for soil B

ESP=(0.1560/10.8485)*100= 1.44%